What is the surface area of the exterior walls of a building with the dimensions of 36' wide x 60' long x 8' height?
Found 2 solutions by stanbon, rothauserc:
Answer by stanbon(75887)

What is the surface area of the exterior walls of a building with the dimensions of 36' wide x 60' long x 8' height?
----------------------------------------------
Ans: Area = 2[36*60+36*8+60*8) = 5856 sq. ft.
-------------------------------
Cheers,
Stan H.
---------------------

Answer by rothauserc(4718)

Perimeter of exterior walls = 2*(60+8) + 2*(36+8) = 224 feet